# Revisiting Graph Analytics Benchmark

### v1

@@ 4       $c\leftarrow 0,\ j\leftarrow i$
9            if $u_{k}$.degree $<u_{k}$.degree_limit then
@@ 4.2.1. Density Enhancement
To enhance the density and generate more edges, $\mathsf{LDBC}$-$\mathsf{DG}$ adds a lower bound on the probability $p_{limit}$. However, this resolution is overly simplistic and lacks justification, as assigning the same connection probability $p_{limit}$ to all distant vertices, regardless of their actual separation.
Density Factor. Our approach is adding a density factor $\alpha>1$ (which is used to control the density of the generated graph) during the sampling process to enhance the generation of edges, as shown in Line 5 of Algorithm 1. The factor $\alpha$ serves to concentrate probability from distant to local vertices. More precisely, we can reformulate the probability of an edge being the first existing edge as follows: $\Pr[e(i,j)]=\frac{1}{c+\frac{d(d-1)}{c}+2d-1}$, where $d=j-i$ is the distance. The problem is the change when $c$ is reduced to $\frac{c}{\alpha}$. An observation is that $\Pr[e(i,j)]$ has a pair of same values when $c=\frac{d(d-1)}{c}$. If $c<\frac{d(d-1)}{c}$, $Pr[e(i,j)]$ is always smaller. Otherwise, only when $\frac{c}{\alpha}\leq\frac{d(d-1)}{c}$ can $\Pr[e(i,j)]$ become lower. Combining two conditions, we have:
@@
For convenience, we can use $d\geq\sqrt{c^{2}/\alpha}$ to quickly check since $d\approx d-1$ when $d$ is large. Therefore, if $d$ is larger than this boundary, i.e., the vertices are too far away, the probability $\Pr[e(i,j)]$ is decreasing due to the influence of $\alpha$. Correspondingly, other nearing vertices will have a higher probability. When the value of $\alpha$ is larger, the probability of each new edge being close to the edges generated in the previous round increases, which leads to a higher graph density. Though it is difficult to quantify the impact of $\alpha$, our experiments show that increasing $\alpha$ ten-fold results in approximately twice as many edges being generated.
4.2.2. Diameter Adjustment.
Figure 3. Generating graphs with adjustable diameters
The adjustment involves an additional check in Line 7 of Algorithm 1 that the iteration also breaks once the vertices with indices $i$ and $k$ do not belong to the same group as:
$\left\lfloor\frac{i}{group\_size}\right\rfloor\neq\left\lfloor\frac{k}{group\_size}\right\rfloor$
Once $k$ is too large to exceed the group, the generation process for $u_{i}$ is finished.
4.3. Synthetic Datasets
Specifically, we design four different scales, where the vertex numbers are correlated to the LDBC default setting, i.e., $3.6$M, $27.2$M, $77$M, and $210$M. We also provide two datasets with an enhanced density (i.e., $1.2$M and $9.1$M vertices and the same number of edges) and two datasets with a larger diameter (i.e., $\sim 100$ with the same scale).
5. Benchmarking API Usability
5.1. An Overlooked Benchmarking Scope
The API usability is a qualitative characteristic that evaluates how easy it is to learn and use (Rauf et al., 2019; Myers and Stylos, 2016). Highly usable APIs enhance productivity and reduce bugs, while initial interactions with poorly designed APIs can leave a lasting impression of complexity, potentially deterring users from embracing the platform. The most common evaluation method is empirical studies, where programmers are recruited to answer interview questions (Farooq and Zirkler, 2010; Piccioni et al., 2013b). However, this approach is costly and not scalable, as recruiting participants with adequate programming skills is challenging, and the number of possible usage scenarios for an API can be vast (Farooq and Zirkler, 2010). Consequently, most graph analytics platform benchmarks overlook API usability evaluations, instead reporting the lines of code required for specific tasks as an indicator of API usability (Zhu et al., 2014; Nam et al., 2019; Watson, 2009). Clearly, this approach is oversimplified and inadequate.
@@ 5.2. A Multi-Level LLM-Based Usability Evaluation Framework
Overview of the Framework. Our idea is to simulate programmers with varying levels of proficiency by using LLMs with multi-level prompts. By analyzing the generated code quality, we can evaluate the usability of APIs provided by different platforms. Specifically, we (1) perform instruction-tuning on LLMs to train the Code Generator for each platform by using their public documentation; (2) use the Code Generator to produce the code we requested by providing multi-level prompts; and (3) employ the instruction-tuned LLM-based Code Evaluator to score the generated code.
@@
Step 1: Instruction-Tuning of LLMs. Step 1 aims to let LLMs generate code aligned with each platform’s style, standards, and best practices, especially regarding platform-specific APIs by learning from publicly available documents. As shown in Figure 4 ①, we gather and preprocess data from each platform, including research papers, API documentation, sample code, and coding guidelines, enabling LLMs to learn specific coding styles and conventions. Human evaluators then review the generated results, providing feedback and prompts, which refine the data and the instruction-tuning process. This iterative process continues until the LLM consistently generates code that complies with platform standards.
Figure 5 illustrates how the LLM is trained to generate PageRank code on a specific platform. The example shows its code’s evolution including from general C++ to platform-specific APIs. The LLM is given instructions and prompts to adjust its understanding of platform-specific functions, refining the generated code until it aligns with the platform’s best practices, such as using functions like vertexMap and edgeMapFunction.
@@
Step 2: Multi-Level Prompts. Step 2 simulates programmers with varying expertise levels: junior, intermediate, senior, and expert, as shown in Figure 4 ②. The Code Generator uses varying prompts to generate corresponding code, and we evaluate difficulty by examining the generated code’s quality, thereby assessing the platform’s API usability. For example, high scores at lower-level prompts indicate ease of use and low learning costs of APIs on this platform. The prompt levels and their respective details are as follows:
@@
Step 3: Code Evaluation. Step 3 is crucial for determining the quality of the generated code and the API’s usability. We first define the evaluation metrics. Correctness and readability are commonly used metrics in API usability evaluation (Rauf et al., 2019; Myers, 2017; Piccioni et al., 2013a). Correct code indicates that the API is intuitive and that the provided documentation and examples effectively guide developers in implementing the desired functionality correctly. High readability suggests that the API’s design promotes clear and concise coding practices. These two metrics are crucial for assessing usability. However, we observe that LLMs often exhibit “hallucination”, focusing on general programming patterns or inventing nonexistent API functions while ignoring platform-specific APIs, typically due to unclear or ambiguous prompts. This limitation mirrors the behavior of human developers: less experienced programmers are more likely to make mistakes when dealing with poorly designed APIs. To address this, we introduce a new metric, compliance, which measures how closely the generated code aligns with standard code. This metric reflects the API’s intuitiveness and accessibility for developers with varying skill levels. By assessing whether the API enables users to easily produce code that establishes best practices and standards, compliance provides an objective basis for scoring API usability.
The details of evaluation metrics and their weight are as follows:
  • Compliance (35%). Checking the adherence to platform-specific coding standards and best practices by comparing the generated code with standard code examples. Compliance ensures that the code integrates well with existing systems
@@
A simple method would be to assign equal weights, but this leads to an indivisibility issue. Considering that compliance and correctness more directly reflect the LLM’s understanding of platform APIs compared to readability, we assign slightly higher weights to compliance and correctness. Users can also customize the weight distribution based on their specific needs. During evaluation, we provide Code Evaluator with the evaluated and standard code and ask it to generate scores for each metric based on our provided guidelines. Figure 4③ illustrates the training process: we first provide detailed scoring criteria as LLM’s knowledge base and set basic requirements instructions to get Code Evaluator. Then, we introduce some test codes and provide feedback based on the output results to optimize the Code Evaluator’s instructions. We iterate this process until the LLM can produce stable and satisfactory evaluation results.
@@ 6. Overall Process of Our Benchmark
In this section, we provide a detailed overview of our benchmark’s workflow, aiming to offer an in-depth evaluation of platforms from both performance and usability. As illustrated in Figure 6, the architecture of our benchmark includes three main stages: data preparation, benchmark testing and result analysis.
@@
In the API usability evaluation stage, the Code Generator uses prompts prepared earlier to produce code, which is then combined with the platform-provided code and fed into the Code Evaluator for scoring. To reduce bias, we repeat this process multiple times and average the results to obtain the final score.
@@ 7.1. Environment
We evaluate seven widely used graph analytics platforms. Each platform employs distinct programming languages, storage formats, communication methods, and computation models, which is summarized in Table 6.
@@
  • GraphX (GX) (Gonzalez et al., 2014) is a component in Spark adapted for graph-parallel computation, which provides $\mathsf{vertex}$-$\mathsf{centric}$  computing model using Spark Resilient Distributed Datasets (RDD) APIs.
  • PowerGraph (PG) (Low et al., 2010) is an $\mathsf{edge}$-$\mathsf{centric}$ distributed graph analysis platform, addressing the difficulties including load imbalance in graph algorithms over real-world power-law graphs.
  • Flash (FL) (Li et al., 2023b) is a distributed platform, supporting various complex graph algorithms including clustering, centrality, traversal, etc. It extends the $\mathsf{vertex}$-$\mathsf{centric}$ model with global vertex status to make it easy to program complex graph algorithms.
  • $\mathsf{Grape}$ (GR) (Fan et al., 2017b) is a $\mathsf{block}$-$\mathsf{centric}$ parallel platform, which is designed to execute existing (textbook) sequential graph algorithms with only minor modifications in the distribution context.
  • Pregel+ (PP) (Yan et al., 2015) extends the original $\mathsf{vertex}$-$\mathsf{centric}$ platform Pregel (Malewicz et al., 2010) by introducing vertex mirroring and efficient message reduction, making it effective for power-law and dense graphs.
  • Ligra (LI) (Shun and Blelloch, 2013b) is a lightweight $\mathsf{vertex}$-$\mathsf{centric}$ processing platform for shared memory on a single machine with multiple cores. Ligra is applicable if one machine can allocate the whole graph.
  • G-thinker (GT) (Yan et al., 2020) is a $\mathsf{subgraph}$-$\mathsf{centric}$ platform specialized for graph mining problems, which can create and schedule subgraphs with arbitrary shapes and sizes.
All experiments are evaluated on a cluster with $16$ machines. Each machine has $4$ Intel® Xeon® Platinum 8163 @ 2.50GHz CPUs, 512 GB memory, and 3 TB disk space. The 16 machines are connected with a 15 Gbps LAN network.
@@
Throughput Appendix (gab, [n. d.]) $\mathsf{PR}$, $\mathsf{SSSP}$, $\mathsf{TC}$ S8-Std, S8-Dense, S8-Diam, S9-Std, S9-Dense, S9-Diam 32 16
Stress Test Appendix (gab, [n. d.]) $\mathsf{PR}$ S8-Std, S9-Std, S9.5-Std, S10-Std 32 16
@@ 7.3. Datasets
We test seven synthetic datasets, as detailed in Section 4.3. Note that evaluations on real-life datasets are similar to the LDBC benchmark (Iosup et al., 2016) and are therefore not the focus of this work. As mentioned above, these datasets are in four scales: S8, S9, S9.5, and S10, with S10 being the largest one. Additionally, there are three variants for S8 and S9: Std, Dense, and Diam. Here, Std represents the standard dataset ($\alpha$=10), Dense denotes a denser dataset ($\alpha$=1000), and Diam indicates a dataset with a larger diameter ($\alpha$=10).
@@ 7.4. Methodology
  • Statistics Impact assesses platform sensitivity to various dataset characteristics by running all algorithms on the three S8 datasets. This experiment highlights platform optimizations that enable effective handling datasets with diverse characteristics, which may not be evident when using a single dataset alone.
@@
  • Throughput measures data processing capacity. We run the default algorithms on $S8$ and $S9$ in $16$ machines, each configured with $32$ threads.
@@
  • Usability Evaluation involves the pioneering subjective assessment of the platform’s API usability using our LLM-based evaluation framework, as described in Section 5.2. This evaluation is powered by the GPT-4o model (GPT, 2024). To ensure the evaluation is convincing, we limit our evaluation to code provided by the platform and the original authors, excluding any code we have written ourselves.
@@ 8.1. Generation Similarity and Efficiency
Generation Similarity. To illustrate the better alignment with real-world graph characteristics, we follow the previous evaluation methodology (Prat-Pérez and Dominguez-Sal, 2014), i.e., generate communities over the social network and present the distribution of community statistics in the dataset. We choose the LiveJournal (Leskovec and Krevl, 2014) dataset as the ground truth and select six statistics, Clustering Coefficient, Triangle Participation Ratio, Bridge Ratio, Diameter, Conductance, and Size. We use the Jensen-Shannon Divergence to evaluate the similarity and the result in Table 8 suggests that our datasets achieve $2$x lower divergence on average, verifying that $\mathsf{FFT}$-$\mathsf{DG}$ can generate social networks that follow the real-world distribution and are similar to real-world datasets. Figure 7 presents the distribution of three major statistics with complete results provided in the Appendix (gab, [n. d.]).
@@
Generator CC TPR BR Diam Cond Size
@@
Generation Efficiency. The efficiency evaluation involves generating multiple datasets with the density factor $\alpha$ varying in $\{1,10,100,1000\}$. Figure 9 illustrates the total number of trials and generated edges per second, respectively. $\mathsf{FFT}$-$\mathsf{DG}$ demonstrates consistent throughput as density increases, requiring about $1.5$ trials per edge, while $\mathsf{LDBC}$-$\mathsf{DG}$ necessitates more than $8$ trials per edge. Although $\mathsf{LDBC}$-$\mathsf{DG}$ exhibits a faster sampling process, capable of executing up to 2 times the number of trials per second, its edge generating speed is still about 2.2 times slower than $\mathsf{FFT}$-$\mathsf{DG}$.

### v2

@@ 4  $c\leftarrow 0,\ j\leftarrow i$
9    if $u_{k}$.degree $u_{k}$.degree_limit then
@@ 4.2.1. Density Enhancement
To enhance the density and generate more edges, $\mathsf{LDBC}$-$\mathsf{DG}$ adds a lower bound on the probability $p_{limit}$. However, this resolution is overly simplistic and lacks justification, as it assigns the same connection probability $p_{limit}$ to all distant vertices, regardless of their actual separation.
Density Factor. Our approach is adding a density factor $\alpha>1$ (which is used to control the density of the generated graph) during the sampling process to enhance the generation of edges, as shown in Line 5 of Algorithm 1. The factor $\alpha$ serves to concentrate probability from distant to local vertices. More precisely, we can reformulate the probability of an edge being the first existing edge as follows: $\Pr[e(i,j)]=\frac{1}{c+\frac{d(d-1)}{c}+2d-1}$, where d=jid=j-i is the distance. The problem is the change occurs when $c$ is reduced to $\frac{c}{\alpha}$. An observation is that $\Pr[e(i,j)]$ has a pair of the same values when $c=\frac{d(d-1)}{c}$. If $c<\frac{d(d-1)}{c}$, $Pr[e(i,j)]$ is always smaller. Otherwise, only when $\frac{c}{\alpha}\leq\frac{d(d-1)}{c}$ can $\Pr[e(i,j)]$ become lower. Combining two conditions, we have:
@@
For convenience, we can use $d\geq\sqrt{c^{2}/\alpha}$ to quickly check since $d\approx d-1$ when $d$ is large. Therefore, if $d$ is larger than this boundary, i.e., the vertices are too far away, the probability $\Pr[e(i,j)]$ decreases due to the influence of $\alpha$. Correspondingly, other nearby vertices will have a higher probability. When the value of $\alpha$ is larger, the probability of each new edge being close to the edges generated in the previous round increases, which leads to a higher graph density. Though it is difficult to quantify the impact of $\alpha$, our experiments show that increasing $\alpha$ ten-fold results in approximately twice as many edges being generated.
@@
4.2.2. Diameter Adjustment.
Figure 3. Generating graphs with adjustable diameters
The adjustment involves an additional check in Line 7 of Algorithm 1 that the iteration also breaks once the vertices with indices $i$ and $k$ do not belong to the same group as: $\left\lfloor\frac{i}{group\_size}\right\rfloor\neq\left\lfloor\frac{k}{group\_size}\right\rfloor$. Once $k$ is too large to exceed the group, the generation process for $u_{i}$ is finished.
4.3. Synthetic Datasets
Specifically, we design four different scales, where the vertex numbers are correlated to the LDBC default setting, i.e., 3.63.6M, 27.227.2M, 7777M, and 210210M. We also provide two datasets with an enhanced density (i.e., 1.21.2M and 9.19.1M vertices and the same number of edges) and two datasets with a larger diameter (i.e., $\sim 100$ with the same scale).
5. Benchmarking API Usability
5.1. An Overlooked Benchmarking Scope
The API usability is a qualitative characteristic that evaluates how easy it is to learn and use (DBLP:journals/csr/RaufTP19; DBLP:journals/cacm/MyersS16). Highly usable APIs enhance productivity and reduce bugs, while initial interactions with poorly designed APIs can leave a lasting impression of complexity, potentially deterring users from embracing the platform. The most common evaluation method is the empirical study, where programmers are recruited to answer interview questions (farooq2010api; piccioni2013empirical). However, this approach is costly and not scalable, as recruiting participants with adequate programming skills is challenging, and the number of possible usage scenarios for an API can be vast (farooq2010api). Consequently, most graph analytics platform benchmarks overlook API usability evaluations, instead reporting the lines of code required for specific tasks as an indicator of API usability (6976096; DBLP:conf/kbse/NamHMMV19; watson2009improving). Clearly, this approach is oversimplified and inadequate.
@@ 5.2. A Multi-Level LLM-Based Usability Evaluation Framework
Overview of the Framework. Our idea is to simulate programmers with varying skill levels by using LLMs with multi-level prompts. By analyzing the quality of the generated code, we can evaluate the usability of platform APIs. Specifically, we (1) perform instruction-tuning on LLMs to train the Code Generator for each platform by using their public documentation; (2) use the Code Generator to produce the code we requested by providing multi-level prompts; and (3) employ the instruction-tuned LLM-based Code Evaluator to score the generated code.
@@
Step 1: Instruction-Tuning of LLMs. Step 1 aims to enable LLMs to generate code aligned with each platform’s style, standards, and best practices, particularly in the use of platform-specific APIs, by learning from publicly available documents. As shown in Figure 4 ①, we gather and preprocess data from each platform, including research papers, API documentation, sample code, and coding guidelines, enabling LLMs to learn specific coding styles and conventions. Human evaluators then review the generated results, providing feedback and prompts, which refine the data and the instruction-tuning process. This iterative process continues until the LLM consistently generates code that complies with platform standards.
Figure 5 illustrates how the LLM is trained to generate PageRank code on a specific platform. The example shows the evolution of the code, from general C++ to platform-specific API usage. The LLM is given instructions and prompts to adjust its understanding of platform-specific functions, refining the generated code until it aligns with the platform’s best practices, such as using functions like vertexMap and edgeMapFunction.
@@
Step 2: Multi-Level Prompts. Step 2 simulates programmers with varying expertise levels: junior, intermediate, senior, and expert, as shown in Figure 4 ②. The Code Generator uses varying prompts to generate corresponding code, and we evaluate the difficulty of using the API by examining the generated code’s quality, thereby assessing the platform’s API usability. For example, high scores at lower-level prompts indicate ease of use and low learning costs of APIs on this platform. The prompt levels and their respective details are as follows:
@@
Step 3: Code Evaluation. Step 3 is crucial for determining the quality of the generated code and the API’s usability. We first define the evaluation metrics. Correctness and readability are commonly used metrics in API usability evaluation (DBLP:journals/csr/RaufTP19; DBLP:conf/icse/Myers17; DBLP:conf/esem/PiccioniFM13). Correct code indicates that the API is intuitive and that the provided documentation and examples effectively guide developers in implementing the desired functionality correctly. High readability suggests that the API’s design promotes clear and concise coding practices. These two metrics are crucial for assessing usability. However, we observe that LLMs often exhibit “hallucination”, focusing on general programming patterns or inventing nonexistent API functions while ignoring platform-specific APIs, typically due to unclear or ambiguous prompts. This limitation mirrors the behavior of human developers: less experienced programmers are more likely to make mistakes when dealing with poorly designed APIs. To address this, we introduce a new metric, compliance, which measures how closely the generated code aligns with standard code. This metric reflects the API’s intuitiveness and accessibility for developers with varying skill levels. By assessing whether the API enables users to easily produce code that establishes best practices and standards, compliance provides an objective basis for scoring API usability. The details of evaluation metrics and their weights are as follows:
  • Compliance (35%). Checking the adherence to platform-specific coding standards and best practices by comparing the generated code with standard code examples. Compliance ensures that the code integrates well with existing systems.
@@
A simple method would be to assign equal weights, but this leads to an indivisibility issue. Considering that compliance and correctness more directly reflect the LLM’s understanding of platform APIs compared to readability, we assign slightly higher weights to compliance and correctness. Users can also customize the weight distribution based on their specific needs. During evaluation, we provide Code Evaluator with the evaluated and standard code and ask it to generate scores for each metric based on our provided guidelines. Figure 4③ illustrates the training process: we first provide detailed scoring criteria as the LLM’s knowledge base and set basic requirements instructions to get Code Evaluator. Then, we introduce some test codes and provide feedback based on the output results to optimize the Code Evaluator’s instructions. We iterate this process until the LLM can produce stable and satisfactory evaluation results.
@@ 6. Overall Process of Our Benchmark
In this section, we provide a detailed overview of our benchmark’s workflow, aiming to offer an in-depth evaluation of platforms from the perspectives of both performance and usability. As illustrated in Figure 6, the architecture of our benchmark includes three main stages: data preparation, benchmark testing, and result analysis.
@@
In the API usability evaluation stage, the Code Generator uses predefined prompts to produce code, which is combined with the platform-provided code and scored by the Code Evaluator. To reduce bias, we repeat this process multiple times and average the results to obtain the final score.
@@ 7.1. Environment
We evaluate seven widely used graph analytics platforms. Each platform employs distinct programming languages, storage formats, communication methods, and computation models, which are summarized in Table 6.
@@
  • GraphX (GX) (gonzalez2014graphx) is a component in Spark adapted for graph-parallel computation, which provides the $\mathsf{vertex}$-$\mathsf{centric}$ model using Spark Resilient Distributed Datasets (RDD) APIs.
  • PowerGraph (PG) (Low+al:uai10graphlab) is an $\mathsf{edge}$-$\mathsf{centric}$ distributed graph analysis platform, addressing the difficulties including load imbalance in graph algorithms over real-world power-law graphs.
  • Flash (FL) (10184838) is a distributed platform, supporting various complex graph algorithms including clustering, centrality, traversal, etc. It extends the $\mathsf{vertex}$-$\mathsf{centric}$ model with global vertex status to make it easy to program complex graph algorithms.
  • $\mathsf{Grape}$ (GR) (10.1145/3035918.3035942) is a $\mathsf{block}$-$\mathsf{centric}$ parallel platform, which is designed to execute existing (textbook) sequential graph algorithms with only minor modifications in the distribution context.
  • Pregel+ (PP) (yan2015effective) extends the original $\mathsf{vertex}$-$\mathsf{centric}$ platform Pregel (malewicz2010pregel) by introducing vertex mirroring and efficient message reduction, making it effective for power-law and dense graphs.
  • Ligra (LI) (DBLP:conf/ppopp/ShunB13) is a lightweight $\mathsf{vertex}$-$\mathsf{centric}$ processing platform for shared memory on a single machine with multiple cores. Ligra is applicable if one machine can allocate the whole graph.
  • G-thinker (GT) (yan2020g) is a $\mathsf{subgraph}$-$\mathsf{centric}$ platform specialized for graph mining problems, which can create and schedule subgraphs with arbitrary shapes and sizes.
All experiments are conducted on a cluster with $16$ machines. Each machine has $4$ Intel® Xeon® Platinum 8163 @ 2.50GHz CPUs, 512 GB memory, and 3 TB disk space. The 16 machines are connected with a 15 Gbps LAN network.
@@
Throughput Appendix (gab) $\mathsf{PR}$, $\mathsf{SSSP}$, $\mathsf{TC}$ S8-Std, S8-Dense, S8-Diam, S9-Std, S9-Dense, S9-Diam 32 16
Stress Test Appendix (gab) $\mathsf{PR}$ S8-Std, S9-Std, S9.5-Std, S10-Std 32 16
@@ 7.3. Datasets
We test seven synthetic datasets, as detailed in Section 4.3. Note that evaluations on real-life datasets are similar to the LDBC benchmark (DBLP:journals/pvldb/IosupHNHPMCCSAT16) and are therefore not the focus of this work. As mentioned above, these datasets are in four scales: S8, S9, S9.5, and S10, with S10 being the largest one. Additionally, there are three variants for S8 and S9: Std, Dense, and Diam. Here, Std represents the standard dataset ($\alpha$=10), Dense denotes a denser dataset ($\alpha$=1000), and Diam indicates a dataset with a larger diameter ($\alpha$=10).
@@ 7.4. Methodology
  • Statistics Impact assesses platform sensitivity to various dataset characteristics by running all algorithms on the three S8 datasets. This experiment highlights platform optimizations that enable effective handling of datasets with diverse characteristics, which may not be evident when using a single dataset alone.
@@
  • Throughput measures data processing capacity. We run the default algorithms on $S8$ and $S9$ across $16$ machines, each configured with $32$ threads.
@@
  • Usability Evaluation involves the pioneering subjective assessment of the platform’s API usability using our LLM-based evaluation framework, as described in Section 5.2. This evaluation is powered by the GPT-4o model (GPT-4o). To ensure the evaluation is convincing, we limit our evaluation to code provided by the platform and the original authors, excluding any code we have written ourselves.
@@ 8.1. Generation Similarity and Efficiency
Generation Similarity. To illustrate better alignment with real-world graph characteristics, we follow the previous evaluation methodology (prat2014community), i.e., generate communities over the social network and present the distribution of community statistics in the dataset. We choose the LiveJournal (snapnets) dataset as the ground truth and select six statistics, Clustering Coefficient, Triangle Participation Ratio, Bridge Ratio, Diameter, Conductance, and Size. We use the Jensen-Shannon Divergence to evaluate the similarity and the result in Table 8 suggests that our datasets achieve 22x lower divergence on average, which verifies that $\mathsf{FFT}$-$\mathsf{DG}$ can generate social networks that follow the real-world distribution and are similar to real-world datasets. Figure 7 presents the distribution of three major statistics with complete results provided in the Appendix (gab).
@@
Generator $C$ TPR BR Diam Cond Size
@@
Generation Efficiency. The efficiency evaluation involves generating multiple datasets with the density factor $\alpha$ varying in $\{1,10,100,1000\}$. Figure 9 illustrates the total number of trials and generated edges per second. $\mathsf{FFT}$-$\mathsf{DG}$ demonstrates consistent throughput as density increases, requiring about $1.5$ trials per edge, while $\mathsf{LDBC}$-$\mathsf{DG}$ necessitates more than $8$ trials per edge. Although $\mathsf{LDBC}$-$\mathsf{DG}$ exhibits a faster sampling process, capable of executing up to 2 times the number of trials per second, its edge generation speed is still about 2.2 times slower than $\mathsf{FFT}$-$\mathsf{DG}$.